Re: Shift report 20030113 00:00-08:00

From: Hironori Ito (hito@rcf.rhic.bnl.gov)
Date: Tue Jan 14 2003 - 18:19:55 EST

  • Next message: Apache: "Shift report 20030114 16:00-24:00"
    Hello.  Today (1/14), Bob and I tried to fix H2 HV problem during a 
    short controled access.  Making a short story short, we fixed 2 
    channels, H2T05 and H2T07 are fixed. But at same time, we managed to 
    completely kill H2T01 and H2T03 HV.  Anyway, we have a opening tomorrow 
    and will fix them.  For more detail, read a BRAHMS E-log.  If you still 
    do not know where is our On-Line log book, it is located at 
    http://pii3.brahms.bnl.gov/cgi-bin/elog.pl
    
    you can view a previous shift by clicking by Prev icon.
    
    
    Hiro
    
    Ian Bearden wrote:
    > Hi Jens Ivar, others...
    > This on--again, off--again problem with H2 smells like a bad connection
    > somewhere.
    > If it affects all channels, you should probably check the start signal to
    > the H2 TDCs.  There is one START which is daisy chained, so it might be
    > worthwhile to check the box where the signal originates as well as the cable
    > itself.  It would not be the first time a bad LIMO gave problems.
    > Also, the drifting HV needs to be checked ASAP, as the PMTs *CANNOT* take
    > more than
    > -1800V.
    > Hope this helps
    > Ian
    >  --------------------------------------------------------------------
    > | I.G. Bearden                                                       |
    > | Niels Bohr Institute Tb 3   email:      bearden@nbi.dk             |
    > | Blegdamsvej 17              phone:      (+45) 35 32 53 23          |
    > | København Ø                   FAX:      (+45) 31 42 10 16          |
    > | Danmark                       Www:      http://www.nbi.dk/~bearden |
    >  --------------------------------------------------------------------
    > 
    > -----Original Message-----
    > From: owner-brahms-l@bnl.gov [mailto:owner-brahms-l@bnl.gov]On Behalf Of
    > Apache
    > Sent: 13. januar 2003 14:49
    > To: brahms-l@bnl.gov; olchansk@bnl.gov; videbaek@bnl.gov
    > Subject: Shift report 20030113 00:00-08:00
    > 
    > 
    > 
    > ------------------------------------------------------------
    > Date: Mon Jan 13 08:48:50 2003, Shift: 00:00-08:00
    > Supervisor: JIJ, with:
    > Log book#13 pages 56 - 57
    > 
    > Shift Summary:
    > ------------------------------------------------------------
    > 1.
    > The machine had low luminosity for deutrons. Most of the time while beam the
    > blue luminosity was below 18e11. At times the background was low, down below
    > 20k/s in the H1 and H2 scalers. There were, however, sudden changes in the
    > background when MCR was steering, sending the H1 and H2 scalers up to 150
    > and 200k/s. In summary, overall better background level but not stable.
    > 
    > 2.
    > Run 7350 - 117k trigger 6
    > Run 7351 -  39k trigger 6 (aborted by IOserver error, incomplete in
    > database)
    > Run 7352 - 0.1k trigger 6
    > 
    > At this point we are at 167k trigger 6 with FS @ 1/4 A-pol and MRS @ 3kG
    > A-pol.
    > 
    > Run 7353 - pedestal run
    > Run 7354 - TPC pedestal run
    > Run 7355 - DVM run
    > 
    > 3.
    > During a significant increase in background the DAQ reported IO Server
    > error. Also cathode 3 on T5 tripped, but came up again without a problem.
    > 
    > Top channels 1, 3, 5 and 7 on H2 have been geetting higher and higher actual
    > voltage the past shifts, while the demand voltage has been kept constant
    > at -1600V. Passing -1800V I turned demand voltage down to -1400V, which put
    > actual voltage around -1600V.
    > 
    > Also, the H2 problem reported during the Saturday night shift has returned
    > for unknown reason. This is a problem that shows up in the time average
    > distribution.
    > 
    > ------------------------------------------------------------
    > 
    > 
    > 
    


    This archive was generated by hypermail 2.1.5 : Tue Jan 14 2003 - 18:13:38 EST